Posts

Showing posts from February, 2026

NPP Projection Report 2/21 Open

To do this evening/ was completed: Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Snow Day went very well. Print has been rewound and projectors cleaned. Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: Fixed the closed caption settings for T3 and T7 so the captions actually correspond to the correct auditorium number on the CaptiView. Sorry about that. Cue sheets updated/checked (every shift): Marquee updated?: Took down Marty Supreme , The Moment , No Other Choice , and The Secret Agent and rearranged the remaining titles. Posters updated?: Cleaning lights on for the night?: Projection cart by entrance?: Additional Notes:

NPP Projection Close 2/20

To do this evening/ was completed: Presentation Issues: A guest in theater 7 requested a closed captions device during the second round (the feature was Good Luck, Have Fun, Don't Die ) and asked for assistance with it since the captions displayed were for the wrong film. I had it connected to 7 prior to this but went back downstairs, reconnected it, and noticed the issue myself. I switched to theater 3 as an experiment and, weirdly, it had the correct captions being transmitted even though Sirat  was playing in this auditorium.  7 was connected to 3's audio and vice versa. I'm not exactly sure how the signals got crossed as I've never seen this happen before. The round continued without any issues once I made the switch. Could be worth looking into if a rewiring is needed in these rooms. Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Air Bud DVD was delivered this evening fo...

Projection Admin 2/20

To do this evening/ was completed: Presentation Issues: Preshow Clips/Playlist: Built the A1 playlists for 2/27 and added Scream 7 playlists for this week. Lots of SE playlists built please see notes about tech/trailer needs below. Babies playlist built and scheduled for Tuesday! - Lillian  Transferred and or built the playlists for Cutting Through Rocks, Make Out With Violence, BLKNews, and Ganja and Hess to T3 (scheduled these shows on the LMS) so we can tech them come Monday, or if there's time this weekend. - Lillian  Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): We are waiting on the Femme Photographers trailer, which will be ready by EOD Monday and can then get added to a few playlists for 2/27. - Lillian Ordering/Receiving (content):  Ingested Horror_V3 preshow, They Will Kill You trailer, Forecasts. - Lillian Private Events: * This evening, we should be receiving a BluRay for AirBud . Check in with MOD a little later....

Projection Admin 2/19

To do this evening/ was completed:  Continued DCPing shorts.  DCP'd Death Tape, Clean Slate, Action Bog, Nun's Heart and a Grieving Heart. All of these have also been ingested to the LMS. Only a few more to go.  Also! Just another friendly reminder that we already have the Dark Room slide on the LMS.  Marty Supreme and The Moment will be leaving us soon.  Deleted Arco.  Presentation Issues: Preshow Clips/Playlist: Tech'd Sirât and the Oscar Animated/Live Action Shorts. Made the OCAP playlist for Sirât Perfect Blue and Ganja and Hess playlists built.  Trailer Updates: Special Events/ AV set up: Reached out to Chenelle and Raekwon to see if we could get an eta on the BluRay for Crazy Stupid Love  Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Submitted a request for a key extension for Good Luck have Fun Don't Die since our keys will expire early Sunday morning.  Outgoing/Shipping (content): All Coach Carter and T...

NPP Projection Report 2/19 Close

To do this evening/ was completed: In addition to Marty and Moment , we also have no more shows of Secret Agent and No Other Choice starting tomorrow, so these posters/marquee letters should get taken down. Presentation Issues:  Preshow Clips/Playlist: No lights down cue in any version of the Sirat playlist. I added for the 2/20 ones since the 2/13 playlist was only being used for the one show tonight. Manually brought them down for tonight's show. Trailer Updates: Special Events/ AV set up: American Cyborg: Steel Warrior went well. I didn't see the motor start cue on the second reel so the audience saw a bit of leader. I checked the print at the bench after taking it off the projector and the motor cue was there, but out of frame for 1.85 so it did not appear on screen. Try and check for this when inspecting, for 1.85 the cue should be right by the second sprocket from the top of the frame. Ordering/Receiving (posters & trailers): Ordering/Receiving (content): DCP...

NPP Projection Close 2/18

To do this evening/ was completed: Inspected Pulp Fiction. Great condition, on house reels. Ghost went well, print packaged up, mic back in the booth. Trouble in Paradise also smooth sailing, mic broken down, mic stand beneath the screen. There is some LMS scheduling for coming week that I meant to get to: Sirat OCAP (no playlist yet) and Oscar Shorts shows that were initially listed as OCAP. If someone could get to those tomorrow. Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: Cue sheets updated/checked (every shift): Yes Marquee updated?: Posters updated?: Cleaning lights on for the night?: Projection cart by entrance?: Additional Notes:

Projection Admin 2/18

To do this evening/ was completed: Sirat keys open up tomorrow!  Presentation Issues: Preshow Clips/Playlist: Adjusted and added all necessary masking/audio cues for Trouble in Paradise,  which looks great. I can see that the DCP for the Singai ad was labeled Shingani (??)...so it was a tad bit confusing, but that is the correct video. Mic tested! I added a stand and neatly put the mic up and off to the side.  Trailer Updates: Special Events/ AV set up:  Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Downloaded and ingested the Satoshi 15min preshow Kat sent the other day. This is for Perfect Blue.  DCP'd and ingested all of Desmond's slides for March.  We received the video files for the Dark Room Shorts, the slide of which we still have on the LMS so let's keep it there. I downloaded all the videos files to a folder labeled DarkRoomShorts_2026 on the preshow/trailers drive. Currently, I have Nun's Beach transcoding to the Share...

NPP Projection Close 2/17

To do this evening/ was completed: Presentation Issues: Preshow Clips/Playlist: Built and scheduled a show playlist for Blithe Spirit  Trailer Updates: Special Events/ AV set up: Set up for trivia at Trees Lounge + placed everything back after the event ended How to Make Millions Before Grandma Dies screening in theater 4 this evening went smoothly Ordering/Receiving (posters & trailers): Girls Like Girls  posters were delivered and stored behind 3/7 Ordering/Receiving (content): A new DCP drive for Hank's Saloon  was delivered and ingested locally to theater 1 +  American Cyborg print was delivered and brought up to booth 1 Outgoing/Shipping (content): Print/video Inspection: Spent the majority of the evening inspecting American Cyborg  and placed the print on house reels Asset tracking doc updated/checked?: Yes Maintenance/cleaning: Cue sheets updated/checked (every shift): Yes + added credits information for Ghost Marquee updated?: Posters updated?: Clean...

NPP Projection Admin 2/17

To do this evening/ was completed: Inspected Snow Day. Good condition. On house reels. Also put Time To Kill on house reels. *American Cyborg print arrived around 4:20 this afternoon.* Printed return label for Dog Day. It's on the cart. Held off on Coach Carter-- it's from Paramount but didn't come with any return labels included. Emailed John. Update: Dog Day and Love Battlefield picked up. Tech'd Grandma, all set for tonight. Trouble in Paradise keys live tomorrow. Presentation Issues: Preshow Clips/Playlist: Cristina sent us an advertisement to add to Trouble in Paradise , which was DCP'd and added to the playlist. There is a live intro, but don't know if we have a slide for it specifically, but put one in as a placeholder.  Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Teched a reel of Ghost with Cristina and noted the volume Asset t...

NPP Projection Close 2/16

To do this evening/ was completed: LMS scheduled for the coming week. Dog Day Afternoon went smoothly. Print packaged up, mic back in the booth. Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: Cue sheets updated/checked (every shift): Yes Marquee updated?: Posters updated?: Cleaning lights on for the night?: Projection cart by entrance?: Additional Notes:

NPP Projection Admin 2/16

To do this evening/ was completed: Presentation Issues:  Preshow Clips/Playlist: Went through this week's trailer lists to see what needed updating and added house lights up cues to the Sirat show playlists Trailer Updates: Special Events/ AV set up: Tech'd  Kubo and the Two Strings  this morning and added a house lights up cue at the start of the credits Ran loops for Dog Day Afternoon  and focused the lenses/added new tape to the port glass windows for the 1.85 framing + tech'd the Broadway trailer and intro slide and added a feature volume cue to the SPL Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Snow Day , Pulp Fiction and, strangely, There Will Be Blood prints were delivered this afternoon.  Snow Day is up in 4/6 and the other two are down in booth 1. I've emailed to see what to do with There Will Be Blood since it is not on either locations' upcoming schedules Outgoing/Shipping (content): Printed return labels for Love Battl...

NPP Projection Admin 2/15

To do this evening/ was completed: Presentation Issues: Preshow Clips/Playlist: Added a feature volume cue to Kubo and the Two Strings but will tech this tomorrow morning and adjust it as needed Trailer Updates: Special Events/ AV set up: Double checked the  Corpse Bride show playlist this morning and raised the feature volume...later removed this content from the LMS/IMS after its first round screening Ran loops in theater 1 this morning for  Love Battlefield/ Sundays on Fire. The intro and screening went smoothly. Projectors were cleaned/exhausts closed after the show Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Placed  Love Battlefield  back on its shipping reels and packaged the canisters for a return shipment (return labels needed for these) Print/video Inspection:  Asset tracking doc updated/checked?: Maintenance/cleaning: Cue sheets updated/checked (every shift): Marquee updated?: Removed Arco and...

NPP Projection Close 2/14

To do this evening/ was completed: Filling in for Lillian this evening...it's been a busy night for the front of house staff which has meant being on top of bringing cleaning lights up at the start of every round.  I am letting MODs know Alec gave me the go ahead to leave at the start of the final rounds since I will be here tomorrow morning (I can remote in from home if need be but everything should run as automated). Presentation Issues: Much of the feature content sounded lower than normal with the mostly sold out shows this evening so during my rounds I brought up the volume when necessary to compensate for the crowds Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning:  Cue sheets updated/checked (every shift): Yes Marquee updated?: Posters updated?: Organized the poster cu...

Projection Admin 2/13

To do this evening/ was completed: Preshow bits we are waiting on - Satoshi 15min, Dog Day 5min.  (The Dog Day 5min preshow is available on Google drive. I'm waiting to see if these finish uploading to the Share folder as they are currently processing). - LH Tech'd Corpse Bride, Love & Friendship, How to Make Millions... - all playlists have been updated. LH Ghost print inspected, on house reels. Ditto Dog Day Afternoon. Presentation Issues: Preshow Clips/Playlist: Built all first run playlists for the week of 2/20 including many SE playlists with the exception of some SE R-Rated features that are downloading or a need a few adjustments still. I ingested all Oscar Shorts DCPs (live action, animated and documentary) though I only made playlists for the animated and live action shorts, since we normally don't show the docs. We shall find out more when the schedule comes out. - LH Trailer Updates: All trailers updated in new playlists.  Special Events/ AV set up: Ordering/...

NPP Projection Close 2/13

To do this evening/ was completed:  Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Posters for Tow were delivered and stored behind 3/7 Ordering/Receiving (content): Blithe Spirit ,  BLKNEWS , and  Ganja & Hess   feature content finished transferring this evening and was ingested to the LMS (tried and failed transferring  BLKNEWS  to theater 4, began transferring Blithe Spirit and Ganja & Hess  to theater 3 before leaving) Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: Cue sheets updated/checked (every shift): Double checked all cues for tomorrow and cross referenced the schedule on the LMS Marquee updated?: Posters updated?: Cleaning lights on for the night?: Yes Projection cart by entrance?: Yes Additional Notes: Amazon package for Alec was delivered and brought to 4/6

NPP Projection Close 2/12

To do this evening/ was completed: Presentation Issues: First round in theater 6 ended without light cues for the credits/end of show triggering as automated due to an IMS connectivity issue. Brought up everything manually and reseated the ethernet cable to the sound rack which got it back online. Second round pre-show and feature played with no issues Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: The Hunger  drag show and screening in theater 2 went smoothly...put back all equipment used for the event and brought the mics up to the booth to charge Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Ganja & Hess  finished transferring through CyberDuck this evening and was ingested to the LMS (update: file error alert for this transfer attempt Oscar Nominated Shorts content was delivered via Deluxe this evening Outgoing/Shipping (content): Matthew picked up Jurassic Park this afternoon Print/video Inspection: Got sidetracked by ...

Projection Admin 2/12

To do this evening/ was completed: Went through the slightly more elaborate lighting cues with Will for the first drag performance this evening before second round. He's got it!  Presentation Issues:  Preshow Clips/Playlist:  Trailer Updates: Special Events/ AV set up: Setup the speakers in T2 with the yamaha mixer and wireless mic mixer for The Hunger drag show this evening. Tested everything.  Ordering/Receiving (posters & trailers): Received posters for The Drama and Ready or Not 2, both of which have been stored/labeled.  Ordering/Receiving (content): Received the DCP of Ganja and Hess, which is downloading to the share folder. It should be done a little after I leave.  Sent our certs for T4 to the DCP contact for Blknws later this month, 2/28.  Ingested Perfect Blue (made sure to ingest the one in Japanese w/ English subs).  Ingested Cutting Through to the LMS Outgoing/Shipping (content): Matthew picked up Jurassic Park and I got Dancing ...

NPP Projection Close 2/11

To do this evening/ was completed: Coach Carter went smoothly, print packed up, mic back in booth. T7 server cleared of certain features per Alec's request. Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: Cue sheets updated/checked (every shift): Yes Marquee updated?: Posters updated?: Cleaning lights on for the night?: Projection cart by entrance?: Additional Notes:

Projection Admin 2/11

To do this evening/ was completed: Staff screenings are out. Had our cinema meeting and we're in pretty good shape with content for this week. Please also see the Weekly Update I sent out earlier. More notes below.  Presentation Issues: Preshow Clips/Playlist: Built/Adjusted or added to lots of SE playlists today and built the A1 playlist for Sirat. Few things just need to be techd. - LH Trailer Updates: Special Events/ AV set up: Coach Carter all set for this evening. I took one more look at the La La Land playlist just to make sure everything's good. - LH Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Currrently downloading the DCP of Cutting Through Rocks to the Share folder via filemail desktop app. - LH Ingested Love & Friendship this morning.  Downloaded and ingested Kink preshow.  Transferred The Hunger to T2.  Made and ingested slides for American Cyborg/Ghost - LH DCP of Spy kids sorted! Ingesting to the LMS right now. LH Hun...

NPP Projection Close 2/10

To do this evening/ was completed: Presentation Issues: Preshow Clips/Playlist: Transferred the PS_DogDayAfternoon_5min to the LMS, theater 1, and updated the show playlist, converted and transferred all drag show audio/slides to the LMS and theater 2 for tomorrow's tech, and built a show playlist for the tech with everything in the correct order (volume is set to 5.0) Built and scheduled a pre-show playlist for Coach Carter  tomorrow evening  Trailer Updates: Special Events/ AV set up:  Dancing Lady intro and screening went well + Alec helped set up the turntable for the DJ set this evening (update: turntable #2 was giving the DJ some trouble as its needle/the wiring/I do not know what had a signal that was lilting at best...audio would go in and out and/or playback would be super quiet...he opted to run his records off the one turntable that was working without any issues) Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (cont...

NPP Projection Admin 2/10

Hey all, this is another note on the cue sheets, separate from the one about La La Land I sent earlier. For Thursday, all shows in T1 and T3 have been swapped . This is updated on the cue sheet and the Screenwriter schedule. Please note that all shows in T1 and the first two rounds in T2 on Thursday have adjusted timing. Features for these shows will be starting right at the ticketed showtime. The ticketed time hasn't changed, but preshow and lights down will be earlier than usual. This is all reflected on the cue sheets and I've scheduled these accordingly. Please do not adjust anything so the timing is clear for the front of house staff. Thanks! -A To do this evening/ was completed: LMS scheduled for coming week sans some SE playlists. Wuthering Heights and Good Luck built. Loops ran, projectors cleaned. Spy Kids DCP still going. *The Hunger drag show DCPs need to get made tonight.* and ingest to T2 as the tech is scheduled for tomorrow. Presentation Issues: Preshow Clips/P...

Projection Close 2/9 Lillian

To do this evening/ was completed: Got the film schedule tonight before Dead Lover , so I was able to get a good start on the cue sheets for this upcoming week. Almost everything is filled out with the exception of some special event shows and the new feature, GLHFDD , which I have a note about below (in ordering/receiving).  Presentation Issues: Preshow Clips/Playlist: Double checked the playlist for Dead Lover. Tested the connection again.  Trailer Updates: Special Events/ AV set up: Will ran a very smooth NoBudge show in T1. Lots of guests, everyone went home happy.  Zoom Q&a went well in T3. lots of questions, quite the participatory audience. I cleaned up the mic (and stand. added it for cristina. when she does shows in T3 try to put a stand in there) and put the speaker back in T2 with its cord.  Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Ingested Trouble in Paradise to the LMS We received the tips slide for The Hunger, along...

NPP Projection Admin 2/9

To do this evening/ was completed: Finished inspecting Time To Kill. Back on shipping reels for now. Old letters down, new brunch and 35mm letters up. Set up speaker and mic in T3, tested the HDMI connection with a zoom meeting, audio and image good to go. Speaker is turned off for the time being. I brought it down from T2 and the power extension cord attached to it is from 4/6 fyi. Get Out ingested to LMS and T1. Dog Day Afternoon slide and trailer ingested to LMS. Spy Kids DCP going in T1. Double-checked NoBudge playlist, all set. Sent Will my tech notes (aspect, sound). Seagate hard drive was delivered, in 4/6. Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: Cue sheets updated/checked (every shift): Marquee updated?: Posters updated?: Cleaning lights on for t...

NPP Projection Close 2/8

To do this evening/ was completed: T3 held steady for the rest of the evening. Continued inspecting A Time To Kill, great condition so far. Super Bowl stream went smoothly in T2. At the beginning of the first quarter there were a few glitches on screen with sound cutting out for a second, but that quickly went away and the rest of the game played smoothly. Trees Lounge stream was a different situation: same issue we've had in the past with cable: the Digital Link input going black then resuming the stream over and over again. I tried the usual tricks but couldn't resolve it. Found a stream of the game online but it was some minutes behind the cable broadcast and I didn't want to clue people in to the delayed stream, so I asked MOD if they could let people know that T2 was open to them so long as they didn't order anything in there. Not sure what verdict was rendered. Get Out DCP finished at the end of the night, will ingest tomorrow. Presentation Issues: Preshow Clips/P...

NPP Projection Admin 2/8

To do this evening/ was completed: Presentation Issues: First round in theater 3 ( The Moment @ 11:00) did not start as scheduled since the server disconnected again. I did not realize this until playback failed to start at the top of the pre-show. All the lights on the Christie and touchscreen were green. I rebooted the projector and loaded the playlist manually. The trailers and feature played without issue. I reseted the PIB cable from the rack between rounds which had no effect on the server reading "unreachable" on the LMS but the subsequent round loaded and played automatically Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Double checked and tech'd the show playlists for the Blazing Saddles  private event in theater 6 and Peter Hujar's Day in theater 4 this morning Ran loops in theater 1 before Jurassic Park + the feature screening went smoothly and the projectors were cleaned/exhausts closed after it ended Ordering/Receiving (posters &...

Projection Close 2/7

To do this evening/ was completed: Manual starts went just fine tonight. T3 stable. Lost connection briefly in T6 but only at the end of One Battle, and it didn't interrupt playback of course. I did the usual the resat the ethernet cable and it's all fine. This happens every now and again.  Presentation Issues: Preshow Clips/Playlist:  Trailer Updates: Special Events/ AV set up: Tested the cable connection in T2 since we've got the Super Bowl in there tomorrow. Friendly reminder- all you'll need to do is hit HDMI 16x9 on the christie, choose Digital 4 on the processor and you're good to go. I'd have the volume no higher than 5.0! Sounded pretty good at 4.5 but I'm not sure how big the crowd will be. You might have to press a button every now and again if the monitor goes into standby mode.  Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Raekwon dropped off the blurays for our upcoming private events, one of which is Get Out, which...

NPP Projection Report 2/7 Open

To do this evening/ was completed: Presentation Issues: Carly told me the first round Marty Supreme  in T5 was supposed to be OCAP but wasn't (a guest told her but only after the screening ended). I checked the schedule and the playlist labeled "OCAP" was on there for the first round, but the CCAP version of the feature was what was in the playlist. I swapped in the OCAP version. Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: 1.33 masking is set for Dead Lover in T3 on Monday. Cue added to the playlist. Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): I deleted a lot of old content off the screen servers and LMS. All preshows and trailers we're not currently using, of course, and that I don't expect we'll need any time soon. I also set up automatic deletion just for advertisements (which is how preshow clips are classified) and trailers, they'll get automatically deleted from scre...

NPP Projection Close 2/6

To do this evening/ was completed: A slower night on the projection side of things, despite the many guests in the building this evening. Lots of manual starts and calls for bringing up and down lights (which should continue into tomorrow with the number of overlaps between rounds)   Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Spent the majority of the evening finishing inspecting  Love Battlefield  for Sundays on Fire. Reels 2-5 did not have any leaders or cues so I added those. A couple of the leaders I used were taken off spare trailers from booth 1 and can be reattached after the screening Asset tracking doc updated/checked?: Maintenance/cleaning: Cue sheets updated/checked (every shift): Yes Marquee updated?: Posters updated?: Cleaning lights on for the night?: Yes Projection cart by entrance?: Ye...

NPP Projection Admin 2/6

To do this evening/ was completed: A1 2/13 playlists built, old content removed from servers. Tech'd Peter Hujar, all good. Tech'd Dead Lover too, all good, but we'll need to adjust masking presets to properly cover 1.33 in T3. Sent out a zoom link to the participants of the Q&A plus Lillian and Cristina. Cable tested in T2 for Super Bowl. Rescheduled T1 and T4 for today and tomorrow. Cue sheets updated. Dog Day Afternoon print arrived, in T1. Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Techd and adjusted the playlist for Blazing Saddles, which will be for the private event on Sunday in T6. - LH Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Ingested the 2D version of Goat , we had the 3D one on the LMS -A Ingested new content for the week (what was available) for preshow Sundays on Fire V4, Dog Day afternoon preshow, forecasts, etc. Waiting on a few other bits n bobs. - LH  Ingested Kubo and the T...

NPP Projection Close 2/5

To do this evening/ was completed: Presentation Issues: Theater 3's server froze during the Q&A and lost connection to the internet, forcing me to reboot the projector as the moderator and guests were finishing up their discussion, and delaying the next and final round by 5 minutes Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Removal of the Eye screening and Q&A in theater 3 went smoothly save for the aforementioned technical issue that kept me from unpausing the show playlist and starting the next round Alec got STB up and running as I was handling the Q&A in theater 3...I left the gear beneath the table cloth for the show on Saturday but disconnected the spotlight board to plug the masking motor back into the outlet (I made sure the table was out of the path of the curtain/tested scope and flat settings and it is functioning normally). Mics and laptop are back in the booths charging for the time being. Ordering/Receiving (posters & trailers)...

Projection Admin 2/5 Lillian (Alec on mid)

To do this evening/ was completed: There is a headset in T5 at seat 306. Please pick this up once the feature is over. Thanks! Presentation Issues: Preshow Clips/Playlist: Took a look at The Moment today, which I brought up to 5.5 as some dialogue can be soft, but the music is loud enough.  We're moving the Superbowl to T2, swapping the 6:45 The Moment and 9:15 Send Help shows to 3. Made this adjustment on the schedule and cue sheet for Sunday. -A Trailer Updates: Special Events/ AV set up: Setup for Spoons today in T1. Tested the wireless mics, laptop and spots. All you need to do is turn everything on, plus the breaker for the spots and bring down the mics from their charging ports once it is time. The mic stands are by the table, also set up.In addition to this, the three mics and chairs are all set for Removal of the Eye in T3, which I got eyes on just to double check. --Made sure they were all set before heading out. Tested mics, picture, sound, house and spot lights. All g...

NPP Projection Close 2/4

To do this evening/ was completed: T3 Q&A went smoothly, mics broken down afterward. NoBudge playlist built. One of the shipping reels Let The Right One In came with is broken so I didn't package that reel up yet. We have a spare in T1 if we want to use that? Put Dancing Lady and Jurassic Park on house reels. Zip file of Make Out With Violence downloaded to 4/6 desktop. Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: Cue sheets updated/checked (every shift): Yes Marquee updated?: Posters updated?: Cleaning lights on for the night?: Projection cart by entrance?: Additional Notes:

Projection Admin 2/4

To do this evening/ was completed: Weekly update is out - please look things over as soon as you can and familiarise yourself with any pertinent info for shows you'll be working. Thanks! NOTE - forgot to mention this in the update, but the private event for Sirat will be at a later date (still holding, though) this month. We will be showing this as a first run anyhow, so it'll be on the server still. - LH  Presentation Issues: Preshow Clips/Playlist: Went through the Frankenhooker playlist/feature just to double check everything. All good. No intro/memo.  Trailer Updates: Special Events/ AV set up: Setup the mics and chairs (off to the side) in T3 for Monk in Pieces tonight. Just remember to turn on the mixer and you can put the mics on the chairs if you'd like. Mics have been tested. Friendly reminder - spots remote/power switch in 3/7 booth. I noticed there wasn't a slide for the Q&A in the playlist, so that's been adjusted. LH The direct line from the mic m...

NPP Projection Close 2/3

To do this evening/ was completed: Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Set up for trivia this afternoon and put back all equipment used following the show in theater 1.  Let the Right One In screening went well. All reels are tails out beneath the rewind bench in booth 1. Both projectors were cleaned after the show and the exhausts are shut for the night. Ordering/Receiving (posters & trailers): Dead Lover  posters were delivered and stored behind 3/7 Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning:  Cue sheets updated/checked (every shift): Yes Marquee updated?: Posters updated?: Cleaning lights on for the night?: Yes Projection cart by entrance?: Yes Additional Notes: Received a 3.5mm attenuation aux cable from Amazon and brought it up to 4/6

NPP Projection Admin 2/3

To do this evening/ was completed: LMS scheduled for coming week with some first run exceptions. Tech'd Frankenhooker, Removal of the Eye, Monk in Pieces, Let the Right One In. Volumes logged, playlists updated. Pillion and GOAT  ingested to LMS. Loops ran, projectors cleaned. Pulp Fiction trailer back on its core. Die Hard packaged up. Asked John to double check return shipping location. Second can of the SoF print arrived, it's in T1. *A1 2/6 playlists need to be updated with the Mubi La Grazia trailer sometime in the next few days.* T6 storage is at maximum capacity, and there isn't much I can remove right now. We are fine until Monday 2/9 when something will have to be removed to make room for Sinners. Attempted to download the Dead Lover DCP but the Aspera link was prompting me to commit to some new transfer client download so I held off for the moment.  Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters...

NPP Projection Close 2/2

To do this evening/ was completed: Sam and Bruce double feature went smoothly. Both prints are packaged up with the exception of reel one of Die Hard, still need to remove Walter's trailer from it. Mics back in booth. Presentation Issues: T3 server went down right as as the T1 intro was about to start. Lillian was still around at this point so she restarted the projector and things were stable after that. Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: Cue sheets updated/checked (every shift): Marquee updated?: Posters updated?: Cleaning lights on for the night?: Projection cart by entrance?: Additional Notes:

Projection Admin 2/2/26

To do this evening/ was completed: Posting now since we had some funky stuff with T3. See Presentation notes below. L Removal of the Eye/Monk in Pieces playlists scheduled and DCPs transferring to T3 for tech tomorrow. - L fyi - Please see additional notes section.  Presentation Issues: Glad I stuck around later since, right before the 35mm shows, we got walkied for T3 as the server went down right when the feature (Send Help, second round) started. I power-cycled, resat cables, and we were fine from there on out. Once the show was over, I pulled logs from the projector and a report from the server for DMS. We're stable in third round now. - LH Preshow Clips/Playlist: Kat sent updated forecasts for 2/6. I updated all first run and special event playlists for that week. - LH Trailer Updates: Special Events/ AV set up: Ran loops and reel 1 of Die Hard with a Vengeance  and Unbreakable with Lillian this morning in preparation for the double feature tonight. Lenses focused and pla...

NPP Projection Close 2/1

To do this evening/ was completed: Inspected the rest of Die Hard, we are good to go. Cue sheets checked. Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: Cue sheets updated/checked (every shift): Yes Marquee updated?: Posters updated?: Cleaning lights on for the night?: Projection cart by entrance?: Additional Notes:

NPP Projection Clsoe 1/1/25

To do this evening/ was completed: Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: The Incredible Shrinking Man screening went smoothly. The hosts used a laptop for the pre-show and held a raffle inside the auditorium after the film. I set up for the Trees Lounge post-screening festivities in advance and got both hosts situated as the crowd exited the theater.  All gear used was placed back in its proper storage and the projectors were cleaned/exhausts closed. Flat lenses have been swapped for scope for tomorrow's double feature so these will need to be refocused. We are receiving a red band  Pulp Fiction  trailer tomorrow morning (presumably flat) that loops will also need to be run for. 9 to 5 concluded its brunch run in theater 2 this morning and the feature content was removed from the server following the show Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Print/vid...

Projection Close 1/31

To do this evening/ was completed: No issues this evening in any screening room. Finished employee training module.  Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Remember - tomorrow morning will require a laptop setup, which Alec mostly setup this morning, so it should be squared away, just need to plug them in. There is no playlist for this, just the laptop presentation and the print! Please refer to the memo.  Ordering/Receiving (posters & trailers): Ordering/Receiving (content): This evening, Matthew dropped off both Die Hard and Unbreakable . Outgoing/Shipping (content): A Few Good Men has not been picked up yet despite scheduling a pickup with Sony earlier this week. Sent a message to those folks in our thread. There's a chance they had a driver go to WB instead. We'll see.  Print/video Inspection: Inspected all of Unbreakable and the first four reels of Die Hard , so it's more than halfway done and can get finished...